Find the gradient vector and its modulus of scalar function (x,y,z) = x6y4z0-z3yx3-89 at (3,-2,1)


1
Expert's answer
2021-02-19T06:08:46-0500

Answer

Function

V="x^6y^4z^0-z^3yx^3-89"

So

Gradient vector

"\nabla V=(i\frac{dV}{dx}+j\frac{dV}{dy}+k\frac{dV}{dz})"

Putting value of function

"\nabla V= i(6x^5y^4-3x^2yz^3) +j(4x^6y^3-z^3x^3) +k(-3z^2yx^3)"


Now at point (3,-2,1)

"\nabla V=23382i+23355j+162k"

Now it's magnitude

"|\nabla V|=\sqrt{(23382) ^2+(23355) ^2+(162) ^2) }\\=33048.5"